use the given matrix to preform the operation 3 R3 and select the correct option?
1 -2 1
1 0 4
1 -2 5

a. 3 -2 1
5 -1 0
1 -2 0

b. 9 -6 0
3 -3 12
1 -2 5

c. 3 -6 3
1 -3 4
3 -2 0

d. 1 -2 1
1 0 4
3 -6 15

Answer by jim_thompson5910(35256)

3 R3 means you multiply everything in row 3 by the number 3

Row 3: 1 -2 5
1*3 = 3
-2*3 = -6
5*3 = 15

the other rows remain the same. So that's why the answer is choice D